The distance from the bisector to a point P on the fourth nodal line in a two-point interference pattern is 7.86 cm. The distance between the sources is 0.644 cm. The distance from the midpoint between the two sources to
point P is 18.4 cm. What is the wavelength of the waves?
1
Expert's answer
2020-06-23T13:00:48-0400

Solution.

"n=4;"

"x_4=7.86cm;"

"L=18.4cm;"

"d=0.644cm;"

"\lambda-?;"

"\dfrac{x_n}{L}=\dfrac{(n-\dfrac{1}{2})\lambda}{d};"


"\lambda=\dfrac{x_nd}{(n-\dfrac{1}{2})L};"


"\lambda=\dfrac{x_4d}{(n-\dfrac{1}{2})L};"

"\lambda=\dfrac{7.86cm\sdot0.644cm}{(4-\dfrac{1}{2})18.4cm}=0.0786cm=7.86\sdot10^{-2}cm."

Answer: "\lambda=7.86\sdot10^{-2}cm."